# Craft: Cross-modal Aligned Features Improve Robustness of Prompt Tuning

### v2

@@ Abstract
Prompt Tuning has emerged as a prominent research paradigm for adapting vision-language models to various downstream tasks. However, recent research indicates that prompt tuning methods often lead to overfitting due to limited training samples. In this paper, we propose a Cross-modal aligned feature tuning (Craft) method to address this issue. Cross-modal alignment is conducted by first selecting anchors from the alternative domain and deriving relative representations of the embeddings for the selected anchors. Optimizing for a feature alignment loss over anchor-aligned text and image modalities creates a more unified text-image common space. Overfitting in prompt tuning also deteriorates model performance on out-of-distribution samples. To further improve the prompt model’s robustness, we propose minimizing Maximum Mean Discrepancy (MMD) over the anchor-aligned feature spaces to mitigate domain shift. The experiment on four different prompt tuning structures consistently shows the improvement of our method, with increases of up to $6.1\%$ in the Base-to-Novel generalization task, $5.8\%$ in the group robustness task, and $2.7\%$ in the out-of-distribution tasks. The code will be available at https://github.com/Jingchensun/Craft.footnotetext: Preprint. Under Review.
@@ 1 Introduction
We attribute this overfitting to the lack of regularization in the latent space. To address these challenges, we propose a cross-modal feature alignment method that leverages relative representations [38, 50, 21] to construct the static anchor and stochastic anchors in the dual model branch. The static anchors are hard samples from each modality and are used as prototypes to match with the stochastic samples in every training batch. While the stochastic anchors are dynamic samples that are stochastic and selected during every training batch, all the image stochastic samples and the text stochastic samples share a common semantic space. We show that combining the static anchors and the stochastic anchors can provide effective regularization for the learnable prompt and improve the generalization ability of the prompt model.
Additionally, prompt overfitting is exacerbated by domain discrepancies in downstream tasks [20, 59]. As existing methods primarily focus on acquiring task-specific knowledge from the source domains [73, 56, 24], they often neglect the domain distribution discrepancies from the source to target domains. To address this issue, distribution matching metrics like Maximum Mean Discrepancy (MMD) [62, 8, 33] can be applied. However, due to scaling issues between image and text domains, applying MMD in multi-modal prompt tuning is non-trivial. Thus we propose an image-text joint distribution mean discrepancy minimization to enhance the cross-domain consistency of optimized prompts. Based on the aligned feature space, we simply use the joint distribution of image features and text embeddings to reduce the discrepancy between them. Our empirical results show that the modified MMD method efficiently mitigates overfitting and improves the model’s out-of-distribution recognition capability.
@@ 3 Method
Incorporating Prompt Tuning Parameters into $\{\theta,\phi\}$. In prompt-tuning methods, certain learnable prompts are prepended to the text prompts onto the text branch, denoted by $\Lambda_{\text{text}}\in\left\{{\lambda}_{\text{text}}^{1},{\lambda}_{\text{text}}^{2},\cdots,{\lambda}_{\text{text}}^{L}\right\}$. Similarly, learnable prompts $\Lambda_{\text{visual}}\in\left\{{\lambda}_{\text{visual}}^{1},{\lambda}_{\text{visual}}^{2},\cdots,{\lambda}_{\text{visual}}^{L}\right\}$, are prepended into the image branch. For more details regarding the process of prepend, refer to MaPLe architecture or Figure 1 in the paper. In an image classification task, the tokenized image embeddings and learnable visual prompts are fed into the image encoder [56], resulting in image features $f_{{\theta}}\left({x,{\Lambda}_{\text{visual}}}\right)$. Similarly, for the text encoder, tokenized class embeddings and learnable text prompts are processed in the text encoder, yielding text embeddings $g_{{\phi}}\left({y},{{\Lambda}_{\text{text}}}\right)$. To maintain simplicity in this paper, the notation $\{\theta,\phi\}$ will be used to represent the parameters that include $\{{\Lambda}_{\text{visual}},{\Lambda}_{\text{text}}\}$ respectively.
@@ 3.1 Anchors selection for feature alignment
Necessity of Static and Stochastic Anchors. Static anchors provide reliable reference points throughout training, ensuring consistency during feature comparisons. Conversely, stochastic anchors facilitate exploration within the feature space of the domain. Introducing variability through changing anchors, via stochastic anchors, allows models to escape local minima [64]. Stochastic anchors can be viewed akin to data augmentation in the embedding space enhancing model robustness against minor perturbations, contributing to more stable and generalized performance [3].
@@
3.2 Accounting for the variations in distribution within samples
Although prompt-tuning strategies enhance the performance of vision language models, they have a tendency to overfit to the samples, as described in Section 1. Overfitting leads to the generation of inaccurate and uncertain predictions by models when dealing with samples that either have different distributions or belong to Novel classes [4]. This paper will analyze two main distributions of image data, the in-domain samples $x_{\text{id}}\sim\mathbb{P}^{\text{id}}_{x}$ and the out-of-domain samples $x_{\text{ood}}\sim\mathbb{P}^{\text{ood}}_{y}$. Unless explicitly specified, $x$ is assumed to be an in-domain sample $x_{\text{id}}$ throughout the paper. While data can be derived from multiple distributions, we will concentrate on analyzing two distributions at a time to ensure simplicity and clarity. Extending multiple distributions using the concept of group robustness is achievable [53].
Applying a parametric function to any of the two distributions, $x_{\text{id}}$ and $x_{\text{ood}}$, and updating the function’s parameters to make $\mathbb{P}^{\text{id}}_{x}$ and $\mathbb{P}^{\text{ood}}_{x}$ indistinguishable from each other has been demonstrated to be effective in the literature [16]. The technique is straightforward: estimate a particular metric of disparity between the distributions and tune the parameters of the function transformation in order to minimize this disparity. Although techniques such as KL divergence, Optimal Transport measures, can be effective in measuring distributional discrepancies, the maximum mean discrepancy (MMD) is widely favored due to its applicability to various types of data, including imaging, and its well-established statistical properties. The rationale to employ MMD arises from the following result as mentioned in[16] (page $728$) and [11] (page $292$).
@@ 3.2 Accounting for the variations in distribution within samples
MMD can calculate distributional discrepancy between $\mathbb{P}^{\text{id}}_{x}$ and $P^{\text{ood}}_{x}$, but multi-modal data, such as vision and language, presents challenges. The problem lies in incorporating the text probability measure $\mathbb{P}_{y}$ to minimize MMD between $\mathbb{P}^{\text{id}}_{x}$ and $P^{\text{ood}}_{x}$. An obvious solution is to utilize a product measure $\mathbb{P}_{x}\times\mathbb{P}_{y}$. The product measure theorem (Theorem $4.4.6$ from [11]) enables the use of product measures because there exists a unique measure on the product of the $\sigma-$algebras $\mathcal{F}_{x}\times\mathcal{F}_{y}$. It reminds us of the fundamental,
@@
Cardinality of the $\sigma-$algebra for the product measure is greater than that of the individual probability measures.
Please refer to page $139$ of [11] for more details. The above Fact 3 significantly affects MMD computation, as empirical estimates require more samples now to approximate expectations. As a solution, we propose utilizing an induced measure to evaluate MMD in lieu of product measures. For the image encoder $f_{\theta}$ and a unit-normed text anchor $a_{y}$, let $a_{y}^{T}f_{\theta}(\cdot)$ be a measurable function from $\Omega_{x},\mathcal{F}_{x}$ to $\Gamma_{x},\mathcal{G}_{x}$. Our induced measure, the anchor-aligned probability measure, is
@@
The anchor-alinged probability measure $\mathbb{P}_{x}^{a_{y}}$ produces anchor aligned-feature vectors and anchor-aligned MMD measure defined as follows,
@@
Here $k(x,\cdot)$ indicates that the kernel has one parameter fixed at $x$ second is free, in particular, with the feature mapping onto RKHS as $\Phi$, we have $\langle\Phi(x),\Phi(y)\rangle=k(x,y)$. In the empirical form of the Maximum Mean Discrepancy (MMD), the expectation is replaced with the sample average calculated over a group. A previous study [72] is of particular relevance because it has demonstrated that calculating the average of samples taken from training batches ( sub-samples) is feasible while maintaining consistency and asymptotic normality criteria. Moreover, [72] can be seen as an extension of the paper [2] where the in-domain and out-domain samples are parameterized separately. For a detailed explanation of the consistency and asymptotic properties of MMD and the necessary assumptions, we recommend the reader to visit [71]. However, it is crucial that the anchor-aligned feature vectors require holder continuity. We demonstrate the validity of the statement below
@@
$\|f_{\theta}(x_{\text{id}})^{T}a_{y}-f_{\theta}(x_{\text{ood}})^{T}a_{y}\|\leq\|f_{\theta}(x_{\text{id}})-f_{\theta}(x_{\text{ood}})\|\|a_{y}\|\leq\alpha$ (7)
The lemma referred to as lemma 5 ensures that the consistency properties and the asymptotic guarantees will remain valid for the feature vectors that are aligned with the anchor.
@@ 4.1 Feature Alignment Improves the Base-to-Novel Generalization
The feature alignment method improves the accuracy of both Base and Novel group classes. In Table 1, for the Base group, the final average results were consistently enhanced by applying $\mathcal{L}_{\text{Aligned}}$ to four prompt tuning structures, with a $2.0$-point gain achieved by MaPLe [56] and a $1.9$-point gain achieved by PromptSRC [24] on average. For some individual datasets, applying our method $\mathcal{L}_{\mathrm{Aligned}}$ to MaPLe improved the results by 6.1 points on StanfordCars while applying it to PromptSRC led to a 4.9-point improvement on FGVC-Aircraft. Additionally, our $\mathcal{L}_{\text{Aligned}}$ loss consistently enhanced performance on novel classes. An average gain of $2.3$ points is yielded by $\mathcal{L}_{\text{Aligned}}$ on MaPLe, while an average improvement of $1.5$ points is resulted on PromptSRC. The most significant gain on individual datasets was $5.5$ points on EuroSAT for PromptSRC. Although some datasets, such as Flowers102 and Food101, show only minimal improvements, we attribute this to the already high recognition accuracy (over $90\%$) in these tasks from the original CLIP model, thus the improved space is not as significant as other tasks (like DTD or EuroSAT). In most datasets, better results are achieved by our method compared to the baseline models.
@@
Table 3: Cross Robustness Evaluation on Waterbird and CelebA Datasets. Here we applied our $\mathcal{L}_{\text{Aligned}}$ on the baseline method. "WG" refers to the Worst-Group accuracy and "Avg" refers to the Average-Group accuracy. The A downward arrow $\downarrow$ indicates a reduction in the accuracy gap for the groups.
@@ 4.5 Aligned Feature Visualization
Figure 5 illustrates the qualitative results of our method on a single sample. In Figure 5 (a), without using the $\mathcal{L}_{\text{Aligned}}$ loss function, an image of an English Marigold was incorrectly classified as Marigold in MaPLe. However, after applying the $\mathcal{L}_{\text{Aligned}}$ loss function, the top-1 probability for this image corresponded to the correct category. Figures 5(b) show similar results.
@@ B.5 Maximum Mean Discrepancy Implementation
We define MMD using two equivalent forms: as a distance between feature means and as an integral probability metric. Given random variables $X$ from distribution $P$ and $Y$ from distribution $Q$, the feature map $\beta$ projects $X$ and $Y$ to a new space $\mathcal{\beta}$. Using the kernel trick, the inner product in $\mathcal{\beta}$ is defined as:
$k(X,Y)=\langle\beta(X),\beta(Y)\rangle_{F}$ (9)
Feature means, or mean embeddings, compute the coordinate-wise means of these projections:
$\mu_{P}(\beta(X))=[E[\beta(X_{1})],\dots,E[\beta(X_{m})]]^{T}$ (10)
The inner product of feature means is then expressed in terms of the kernel function:
$\langle\mu_{P}(\beta(X)),\mu_{Q}(\beta(Y))\rangle_{F}=E_{P,Q}[k(X,Y)]$ (11)
The squared MMD value, estimating the distributional discrepancy, is computed as:
$MMD^{2}(P,Q)=\|\mu_{P}-\mu_{Q}\|_{F}^{2}$ (12)
In practical applications, for instance, aligning image features distribution from the source domain to the target domain, MMD helps minimize the discrepancy between cross-domain distributions. This is critical in tasks like domain adaptation, where source domain logit $s^{s}=\operatorname{sim}(u,v)$ and target domain logit $s^{t}=\operatorname{sim}(u^{t},v)$ are matched.
@@ C.2 Group Robustness Baseline
We evaluate our method against several methods in group robustness experiments, including zero-shot classification, ERM linear probing [28], and ERM adapter training [14]. Additionally, we compare against recent approaches tailored to enhancing downstream transfer in analogous scenarios, all while utilizing only pretrained model embeddings.

### v3

@@ Abstract
Prompt Tuning has emerged as a prominent research paradigm for adapting vision-language models to various downstream tasks. However, recent research indicates that prompt tuning methods often lead to overfitting due to limited training samples. In this paper, we propose a Cross-modal Aligned Feature Tuning (CRAFT) method to address this issue. Cross-modal alignment is conducted by first selecting anchors from the alternative domain and deriving relative representations of the embeddings for the selected anchors. Optimizing for a feature alignment loss over anchor-aligned text and image modalities creates a more unified text-image common space. Overfitting in prompt tuning also deteriorates model performance on out-of-distribution samples. To further improve the prompt model’s robustness, we propose minimizing Maximum Mean Discrepancy (MMD) over the anchor-aligned feature spaces to mitigate domain shift. The experiment on four different prompt tuning structures consistently shows the improvement of our method, with increases of up to $6.1\%$ in the Base-to-Novel generalization task, $5.8\%$ in the group robustness task, and $2.7\%$ in the out-of-distribution tasks. The code is available at https://github.com/Jingchensun/Craft.
@@ 1 Introduction
We attribute this overfitting to the lack of regularization in the latent space. To address these challenges, we propose a cross-modal feature alignment method that leverages relative representations [37, 49, 21] to construct the static anchor and stochastic anchors in the dual model branch. While the stochastic anchors are dynamic samples that are stochastic and selected during every training batch, all the image stochastic samples and the text stochastic samples share a common semantic space. We show that combining the static anchors and the stochastic anchors can provide effective regularization for the learnable prompt and improve the generalization ability of the prompt model.
Additionally, prompt overfitting is exacerbated by domain discrepancies in downstream tasks [20, 58]. As existing methods primarily focus on acquiring task-specific knowledge from the source domains [70, 55, 24], they often neglect the domain distribution discrepancies from the source to target domains. To address this issue, we proposed an image-text joint distribution mean discrepancy minimization to enhance the cross-domain consistency of optimized prompts. Our empirical results show that the modified MMD method efficiently mitigates overfitting and improves the model’s out-of-distribution recognition capability.
@@ 3 Method
Incorporating Prompt Tuning Parameters into $\{\theta,\phi\}$. In prompt-tuning methods, certain learnable prompts are prepended to the text prompts onto the text branch, denoted by $\Lambda_{\text{text}}\in\left\{{\lambda}_{\text{text}}^{1},{\lambda}_{\text{text}}^{2},\cdots,{\lambda}_{\text{text}}^{L}\right\}$. Similarly, learnable prompts $\Lambda_{\text{visual}}\in\left\{{\lambda}_{\text{visual}}^{1},{\lambda}_{\text{visual}}^{2},\cdots,{\lambda}_{\text{visual}}^{L}\right\}$, are prepended into the image branch. For more details regarding the process of prepend, refer to MaPLe [55] architecture or Figure 1 in the paper. In an image classification task, the tokenized image embeddings and learnable visual prompts are fed into the image encoder [55], resulting in image features $f_{{\theta}}\left({x,{\Lambda}_{\text{visual}}}\right)$. Similarly, for the text encoder, tokenized class embeddings and learnable text prompts are processed in the text encoder, yielding text embeddings $g_{{\phi}}\left({y},{{\Lambda}_{\text{text}}}\right)$. To maintain simplicity in this paper, the notation $\{\theta,\phi\}$ will be used to represent the parameters that include $\{{\Lambda}_{\text{visual}},{\Lambda}_{\text{text}}\}$ respectively.
@@ 3.1 Anchors selection for feature alignment
Necessity of Static and Stochastic Anchors. Static anchors provide reliable reference points throughout training, ensuring consistency during feature comparisons. Stochastic anchors serve as a form of data augmentation. In each iteration, samples from various modalities are presented to the model, allowing for sampling from the complete support of the data distribution. This approach enhances robustness against minor perturbations and outliers, resulting in more stable and generalized performance. [4].
@@
3.2 Accounting for the variations in distribution
Although prompt-tuning strategies enhance the performance of vision language models, they have a tendency to overfit the samples, as described in Section 1. Overfitting leads to the generation of inaccurate and uncertain predictions by models when dealing with samples that either have different distributions or belong to Novel classes [5]. This paper will analyze two main distributions of image data, the in-domain samples $x_{\text{id}}\sim\mathbb{P}^{\text{id}}_{x}$ and the out-of-domain samples $x_{\text{ood}}\sim\mathbb{P}^{\text{ood}}_{x}$. Extending multiple distributions using the concept of group robustness is achievable [52].
Applying a parametric function to any of the two distributions, $x_{\text{id}}$ and $x_{\text{ood}}$, and updating the function’s parameters to make $\mathbb{P}^{\text{id}}_{x}$ and $\mathbb{P}^{\text{ood}}_{x}$ indistinguishable from each other has been demonstrated to be effective in the literature [16]. Although techniques such as KL divergence, Optimal Transport measures, can be effective in measuring distributional discrepancies, the maximum mean discrepancy (MMD) is widely favored due to its applicability to various types of data, including imaging, and its well-established statistical properties [16] (page $728$) and [11] (page $292$).
@@ 3.2 Accounting for the variations in distribution
MMD measures the distributional discrepancy between in-domain and out-of-domain data; however, challenges occur with multi-modal data, involving vision and language. The straightforward use of MMD necessitates optimization across vision-language, language-language, and vision-vision distribution pairs. We propose calculating the Maximum Mean Discrepancy (MMD) through a vision-language similarity measure. An obvious solution is to utilize a product measure $\mathbb{P}_{x}\times\mathbb{P}_{y}$. The product measure theorem (Theorem $4.4.6$ from [11]) enables the use of product measures because there exists a unique measure on the product of the $\sigma-$algebras $\mathcal{F}_{x}\times\mathcal{F}_{y}$. It reminds us of the fundamental [11],
@@
The cardinality of the $\sigma-$algebra for the product measure is greater than that of the individual probability measures.
Please refer to page $139$ of [11] for more details. The above Fact 3 significantly affects MMD computation, as empirical estimates require more samples now to approximate expectation calculations. As a solution, we propose utilizing an induced measure to evaluate MMD in lieu of evaluating it over product measures. For the image encoder $f_{\theta}$ and a unit-normed text anchor $a_{y}$, let $a_{y}^{T}f_{\theta}(\cdot)$ be a measurable function from $\Omega_{x},\mathcal{F}_{x}$ to $\Gamma_{x},\mathcal{G}_{x}$. Our induced measure, the anchor-aligned probability measure, is
@@
Based on the the anchor-aligned probability measure $\mathbb{P}_{x}^{a_{y}}$ defined in Equation 5, the anchor-aligned MMD measure defined as follows:
@@
Equation 6 defines the MMD loss between two distributions $\mathbb{P}^{\text{id}}_{x}$ (in-domain) and $\mathbb{P}^{\text{ood}}_{x}$ (out-of-domain) over the anchor-aligned feature space. Here, $k(x,\cdot)$ signifies that the kernel has one parameter fixed at $x$, while the second parameter, represented by $\cdot$, is free and can accommodate any random variable. In particular, with the feature mapping onto RKHS as $\Phi$, we have $\langle\Phi(x),\Phi(y)\rangle=k(x,y)$. In the empirical form of the MMD, the expectation is replaced with the sample average calculated over a group. A previous study [69] is of particular relevance because it has demonstrated that calculating the average of samples taken from training batches ( sub-samples) is feasible while maintaining consistency and asymptotic normality criteria. Moreover, [69] can be seen as an extension of the paper [3] where the in-domain and out-domain samples are parameterized separately. For a detailed explanation of the consistency and asymptotic properties of MMD and the necessary assumptions, we recommend the reader to visit [68]. However, it is crucial that the anchor-aligned feature vectors require holder continuity. We demonstrate the validity of the statement below
@@
$\|f_{\theta}(x_{\text{id}})^{T}a_{y}-f_{\theta}(x_{\text{ood}})^{T}a_{y}\|$ $\leq\|f_{\theta}(x_{\text{id}})-f_{\theta}(x_{\text{ood}})\|\|a_{y}\|$
$\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|x_{\text{id}}-x_{\text{ood}}\|^{\alpha}}$ (7)
The first inequality in Equation (7) is valid by the Cauchy-Schwarz inequality. The second inequality is valid because of the unit norm constraint on the anchor aligned vectors $a_{y}$. The lemma 5 ensures that the consistency properties and the asymptotic guarantees will remain valid for the feature vectors that are aligned with the anchor.
@@ 4.1 Feature Alignment Improves the Base-to-Novel Generalization
The feature alignment method improves the accuracy of both Base and Novel group classes. In Table 1, for the Base group, the final average results were consistently enhanced by applying $\mathcal{L}_{\text{Aligned}}$ to four prompt tuning structures, with a $2.0$-point gain achieved by MaPLe [55] and a $1.9$-point gain achieved by PromptSRC [24] on average. For some individual datasets, applying our method $\mathcal{L}_{\mathrm{Aligned}}$ to MaPLe improved the results by 6.1 points on StanfordCars. Additionally, our $\mathcal{L}_{\text{Aligned}}$ loss consistently enhanced performance on novel classes. An average gain of $2.3$ points is yielded by $\mathcal{L}_{\text{Aligned}}$ on MaPLe, while an average improvement of $1.5$ points is resulted on PromptSRC. Although some datasets, such as Flowers102 and Food101, show only minimal improvements, we attribute this to the already high recognition accuracy (over $90\%$) in these tasks from the original CLIP model, thus the improved space is not as significant as other tasks (like DTD or EuroSAT). In most datasets, better results are achieved by our method compared to the baseline models.
@@
Table 3: Cross Robustness Evaluation on Waterbird and CelebA Datasets. Here we applied our $\mathcal{L}_{\text{Aligned}}$ on the baseline method. "WG" refers to the Worst-Group accuracy and "Avg" refers to the Average-Group accuracy. The A downward arrow $\downarrow$ indicates a reduction in the accuracy gap for the groups. The decrease in average accuracy is normal in the group robustness setting, as the goal is to reduce the accuracy gap between groups.
@@ 4.5 Aligned Feature Visualization
Figure 5 illustrates the qualitative results of our method on a single sample. In Figure 5 (a), without using the $\mathcal{L}_{\text{Aligned}}$ loss function, an image of an Tabby cat was incorrectly classified as Egyptian Mau. However, after applying the $\mathcal{L}_{\text{Aligned}}$ loss function, the top-1 probability for this image corresponded to the correct category. Figures 5(b), (c) and (d) show similar results.
@@ B.5 Maximum Mean Discrepancy Implementation
To compute MMD for multi-modal data, a product measure is constructed to create a new probability space. Combining two probability spaces increases the complexity of the resulting $\sigma$-algebra, necessitating additional samples to characterize the probability space, as noted in Fact 3. We propose Equation 5 to define an induced measure, specifically the anchor-aligned probability measure, as a replacement for the traditional product measure in MMD computation. Equation 5 is essential for the application of MMD in anchor-aligned feature spaces. The transformation of the original probability measure $P_{x}$ via an anchor-aligned mapping is demonstrated. This equation defines a new probability measure, $P_{x}^{a_{y}}$, corresponding to the anchor $a_{y}$.
Equation 6 specifies the MMD loss between two distributions, $\mathbb{P}_{{x}}^{id}$ (in-domain) and $\mathbb{P}_{{x}}^{ood}$ (out-of-domain), within the anchor-aligned feature space. This equation quantifies the discrepancy between two probability distributions in the anchor-aligned feature space. Here is how to use Equation 6 for the current task: The first term computes the expectation over all $x_{\text{id }}$ samples, while the second term computes the expectation over all $x_{\text{ood }}$ samples. Here, $k$ is the kernel function, and in our experiments, we use the Gaussian kernel. $f$ is the image encoder, and $\theta$ is being updated during training. In practical implementations, the expectations are replaced with sample averages.
@@ C.2 Group Robustness Baseline
We evaluate our method against several methods in group robustness experiments, including zero-shot classification, ERM linear probing [28], and ERM adapter training [14]. Additionally, we compare against recent approaches tailored to enhancing downstream transfer in analogous scenarios, all while utilizing only pretrained model embeddings [1].